You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130218 (28 Feb 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 February 17-18

Observer: E. Sani, M. Fumana
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC Bin

Summary

Thin cirrus above the South orizion at the beginning of the night, with low humidity and quite strong wing gusts ~15m/s.
Due to the high wind gusts at 4.00UT we have to close the dome.
The sky is clear when we open again the dome.
During the execution of the Grazian program we had an trackers error from the red side. The TO fixed it and the SA re-initialized the software.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard SA101 B-Bessel, R-Bessel, U-Bessel, V-Bessel   airmas 1.2
Saracco_DDT Lynx B-Bessel, R-Bessel 1h COMPLETE
Standard SA104 U-special, r-Sloan   airmass 1.6
Grazian_30A GOODS-N U-special, r-Sloan 2h sky conditions and seeing within the specs
Standard SA101 U-special, r-Sloan   airmass 1.6
Standard SA109 U-special, r-Sloan   airmass 1.4

Details

01:19 The TO opens the dome

01:43 dofpia,/first,/X2
  • failed: background too high

01:46 Try without X2: dofpia,/first
  • failed: background too high. Wait few minutes

01:53 dofpia,/first
  • Airmass ~1.4

02:07 Slew to Saracco
  • pointing ob
  • dofpia,/X2

02:15 lbcrangebal B=021540 R=021537

02:00 Run SA98_BRUV.ob

02.23 Run Lynx_BR_1.ob
  • The average seeing in the blue channel is 1.7 and 1.4 in the red channel.

03.00 Error from the LBC panel. We abort the script. The TO fix it and the SA re-inizialize the software. The 6th blue frame was loosen.

03.03 dofpia,/X2 on target field Saracco

03.12 Run Lynx_BR_1a.ob
  • We edit a new ob with the last 5 random offsets of the original pattern.
  • Thin cirrus passing

03:53 SWITCH TO LUCI

03:55 CLOSING DOME
  • Wind too high. We have to wait until the wind goes down

07:02 Acquiring Dark frames for LUCI

07:23 The TO opens the Done
  • We ask the TO to switch back to LBC

07:55 Dome opened, and we are authorized to use LBC.
  • We slew to the standard star for Grazian_30A

07:58 dofpia,/X2 on the field ofSA101

08:05 Run SA101_Sr.ob

08:14 dofpia,/first,/x2 on the scientific field Grazian
  • We use the /first because even if the previous dofpia converged, the extra-focus pupils did not show the central hole.

08:23 lbcrangebal B=082411 R=082407

08:27 Run grazian_GOODSN_UR_11.ob
  • The average seeing is 1.5

09:07 dofpia,/x2 on the scientific field Grazian

09:13 Run grazian_GOODSN_UR_12.ob
  • We have variable seeing from one frame to another, ranging between 1.1 and 1.7 in both channels.

09:54 dofpia,/x2 on the scientific field Grazian

09:59 Run grazian_GOODSN_UR_13.ob

10:18 We have a problem with the red trackers. We abort the ob.

10:22 The SA fix the problem *dofpia,/X2 on the science filed Grazian

10:30 Restart grazian_GOODSN_UR_13.ob
  • The average seeing is within the specs fro both channels.

11.11 dofpia,/X2 on the science field Grazian

11:19 Run grazian_GOODSN_UR_14.ob

12:00 dofpia,/X2 on the science field Grazian

12.05 Run grazian_GOODSN_UR_15.ob
  • 12.20-12.30 there is a strong jump in the seeing towards bad bad values.

12.50 dofpia,/X2 on the standard field

12:56 Run SA104_Sr.ob
  • airmass 1.6

13:05 We slew to a standard with lower airmass
  • dofpia,/X2

13:12 Run S109_Sr.ob
  • airmass 1.3

13.17 The TO closes the dome.

-- EleanoraSani - 18 Feb 2013
Topic revision: r12 - 28 Feb 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback